OpenStudy (sbentleyaz):

Find the x-coordinates of any relative extrema and inflection point(s) for the function f(x)=9x^1/3 +9/2 x^4/3 . You must justify your answer using an analysis of f '(x) and f "(x)

OpenStudy (mrs.ambrose614):

I'm solving it now bear with me its going to take a while @sbentleyaz

OpenStudy (sbentleyaz):

@3mar

OpenStudy (3mar):

@triciaal If you could help here, I would be grateful!

OpenStudy (sbentleyaz):

@triciaal could you please help

zepdrix (zepdrix):

\[\large\rm f(x)=9x^{1/3} +\frac92x^{4/3} \]What do you need help with? Were you able to find the first derivative function?

OpenStudy (3mar):

@sbentleyaz Please, pay attention here!

OpenStudy (sbentleyaz):

is this correct so far we want to find f'(x) - the first derivative and f"(x)- the second derivative f(x)= 9x^1/3+ 9/2x^4/3 f'(x)=3x^-2/3 + 6x^1/3 f''(x)=-2x^-5/3+2x^-2/3 f'(x) is used to find the local extrema and f''(x) is used to find the inflection points set f'(x) equal to zero and solve. We get x=0 and x=-1/2. We can not have 0 though, because it is in the denominator and you can not divide by 0. therefore the x value of the relative extrema is x=-1/2. substitute this into the original equation and you get Relative Extrema: (-1/2, -5.3) for the inflection points we set f"(x) equal to zero and solve x=0 and x=-1 Again, x can not equal 0 due to it being in the denominator, so x=-1 we plug -1 into the original equation to find the y coordinate so

OpenStudy (sbentleyaz):

this is what i have been working on

zepdrix (zepdrix):

They only wanted the x-coordinate of your relative extrema. So you did a bit more work than they asked for hehe :)

zepdrix (zepdrix):

I think it's positive 1 for the inflection point. Lemme check my work again.

OpenStudy (sbentleyaz):

Okay thank you

zepdrix (zepdrix):

I think x=0 can be an inflection point actually...\[\large\rm 0=-2x^{-5/3}+2x^{-2/3}\]Factoring gives us,\[\large\rm 0=-2x^{-5/3}(1-x)\] We have possible inflection points where the second derivative is `zero` or `undefined`. We should do a line test to see which ones are actually inflection points.

OpenStudy (sbentleyaz):

so inflection at x=0 and x=-1

OpenStudy (sbentleyaz):

and for the relative extrema we do not have one at x=0 correct @zepdrix

zepdrix (zepdrix):

Plugging in test points, \(\large\rm f''(-1)=-2(-1)^{-5/3}(1--1)=positive\) \(\large\rm f''(1/2)=-2\left(\frac12\right)^{-5/3}\left(1-\frac12\right)=negative\) \(\large\rm f''(2)=-2(2)^{-5/3}(1-2)=positive\)

OpenStudy (sbentleyaz):

but the neative fraction exponent will create a denominator where 0 can not be subbed in

OpenStudy (sbentleyaz):

so how could 0 be an answer

zepdrix (zepdrix):